15 April 2014 Lunar Eclipse observations with MODS

slitmask given ID599301 (though there is no ID cut into the mask). MASKNAME = ID599301 in data headers.

Data files are: mods1b.20140415.NNNN.fits

Calibrations

01:43 Start Calibrations: MODS Blue-only, 2x2 binning. Tel @ zenith, lights off in enclosure.
    • Slitflats using mods_moon1 slit (ID599301)
    • 17,18,19 QTH1 + ND1.5 1-sec exposure --- 40,000 counts max
    • 20 QTH1+QTH2 + UG5 0.5-sec(?) exposure --- saturated

01:47 Opening enclosure. Will take further calibrations (comparison lamp spectra + slitless flats, using standard scripts, tomorrow).

Start of Eclipse Observations

05:45 Handover to configure MODS

05:56 Pointing check with MODS

05:57 Sending ACTIVE preset with MODS to collimate (star is 20 deg from moon and at same elevation)

06:03 While collimating for MODS, run dofpia.
  • DX needs manual coma corrections: Z7 = -1000 nm, Z8 = +1500nm.
  • LBC-Red collimated. MODS1 collimated.

S1 UT 06:26 - 07:04

06:09 set S1 in NSIGUI. acqMODS S1.acq
  • g_sdss 10-sec (mods1b.20140415.0021.fits) --- saturated and 22 is with mask at stow.
  • manually set exptime to 1-sec.
  • manually set exptime to 0.5-sec and binning to 1x1 for acq image. better, but still saturated at lower left (24)
  • 25 u_sdss, 0.5-sec. --- too faint, g and 0.5-sec is better.
  • 28 is through-slit image with 15" -Y offset... again g_sdss with 0.5-sec exposure time. 25k counts in moon. 14k in sky.
  • execMODS moonS1.obs 60-sec exposure at this slit position, with blue-only grating and 2x2 binning.
    • 29 saturated on moon.
    • 30 dropped exptime to 30-sec. This gave about 50k counts.
    • Dropped exptime to 15-sec and running a series of 10 exposures. (Will need to monitor).
  • John is attempting LBC-Red exposures --- we must use the NB filter F970N20 filter because the others saturate at the shorted exptimes (0.3 sec).
    • Note - the slit position is not exactly as Florian drew. The slit - running up-down along the X-center - does not seem to cross any maria, however. As I wanted to get started, I did not make any left/right offsets, only an offset of -Y = 15" to move the lunar rim off of the 2nd slit from the top. I will be more careful for future positions, especially S2, if necessary.
    • 31-40 use 15-sec exp time and dimmed from 20k to 10k counts. 06:29 - 06:42 UT
    • 41-46 used 30-sec exptime. Dimmed from 23-18k mean. 06:43-06:53 UT (started with execMODS and not execMODS -e, so it went through the imcslock)
    • 47-52 uses 45-sec exptime. 06:56- (same - did not use execMODS -e --- will do so for subsequent series).
      • Note: 47 shows the effect from earlier saturaton - the "blue daggers" --- see column 1564 at Y=0 - 330.
      • Note: There may be some drift. I see the limb moving down in the 3rd segment.

S2 UT 07:21 - 08:23

07:09 Setting up for S2
  • acqMODS to take 10-sec g_sdss exposure 2x2 binning.
    • 53 is with slit in
    • 54 will be with acquisition image with slit out. Position looks OK as far as sky/moon ratio. But it seems to be rotated with respect to the image from Florian.
    • I tried to change the position angle, which was PA=23.5 (54), by giving a -23.5 offset (55), but it appeared to need still more. Conscious that the time was passing, and noting that the reflectivity for this position vs that drawn is similar, I opted just to go back to the original position, where I took another acquisition image (56), put the slit in and took a test spectroscopic exposure (57). This indicated that 200-sec would be a good exposure time (as predicted), so I ran the script.

07:24 Starting series of 200-sec exposures on S2
    • 58-62 --- The limb has been drifting upwards.
      • While 63 is reading out, I sent an offset along the slit of dY = -23 arcsec.
      • Between 67 and 68, I sent another offset down along the slit, dY = -20 arcsec.
      • 68-69... the red end of the spectrum is getting brighter, counts exceeding 47k (68) and 57k (69).
      • offset along the slit by dY = -5 arcsec and send a script with exptime=100s. (70) and (71).

S3 UT 08:38 - 09:00

08:24 Moving to S3.
  • acqMODS moonS3.acq -- image number 73. We examined it - although the rotation may be off a bit, it looks like the slit is going through good regions.
  • execMODS moonS3_single.obs at 45-sec. 20k counts. Increasing exposure time to 70-sec and starting series (will need to stop and adjust expsoure time as we move out of totality).
  • 75: exptime = 70 sec and giving ~35000 counts at red end.
  • change exptime to 60-sec and run execMODS -f BlueGrating to avoid instconfig and save some time.
    • Note "blue daggers" are no longer present.
  • After 78, change exptime to 45-sec.
  • One more change, to 30-sec.

09:00 The moon is brightening quickly and we are seeing a solar spectrum. Stopping.

09:03 Handover back to LUCI.

Calibrations

12:46 Start MODS calibrations:
  • more through-slit flats:
    • 88-90 VFLAT10 + UG5
    • 91-93 QTH1 + ND1.5
    • 94-96 VFLAT10 + UG5
  • standard calibrations but with 1x1 binning ... repeat for 2x2 binning
  • comparison lamp spectra
    • 110 Hg[Ar]
    • 111 Xe+Kr
    • 112 Ar
  • slitless flats
    • 113-117 VFLAT 5 + Clear
    • 118-122 VFLAT 5 + UG5
  • Bias 3Kx3K 1x1 binning (for acquisition images)
  • Bias 8Kx3K 2x2 binning (for spectra).

-- %USERSIG{OlgaKuhn - 2014-04-15}%

Comments

 
I Attachment Action Size Date Who Comment
acqimage.jpgjpg acqimage.jpg manage 19 K 15 Apr 2014 - 06:37 OlgaKuhn acq image at ~06:20 UT 0.5-sec g_sdss 1x1 binning
Topic revision: r3 - 15 Apr 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback